OpenStudy (anonymous):

how would you factor this ? 20x-8y+12?

OpenStudy (anonymous):

each term has a common factor of \(4\) so "factor it out" that is the best you can do

OpenStudy (anonymous):

i.e. put a 4 outside of some parentheses \[4(something)\]

OpenStudy (anonymous):

so it would be 4(5x-2y+3)?

OpenStudy (anonymous):

why yes

OpenStudy (anonymous):

hell yes omg :D im so proud of myself

OpenStudy (anonymous):

yeah by the end of the evening you will be an math whiz

OpenStudy (anonymous):

i hope and its night over here its 12:14 am lol

OpenStudy (anonymous):

ok by 3 a.m. then

OpenStudy (anonymous):

it is night over here too past my bedtime in fact

OpenStudy (anonymous):

lol i hope i have to pas this practise thing //:

OpenStudy (anonymous):

aw okay good night then :)

OpenStudy (anonymous):

i just said it was past my bedtime not going

OpenStudy (anonymous):

lol yay cause sadly i still need your help :P sorry

OpenStudy (anonymous):

once @satellite73 passes his bed-time, he can't fall asleep.

OpenStudy (anonymous):

looks like you are doing fine on your own what is next?

OpenStudy (anonymous):

(x-10)(x+4)=0 @satellite73 idk how to do this //;

OpenStudy (anonymous):

set each factor equal to zero and solve

OpenStudy (anonymous):

for the first one you get \(x-10=0\) and so \(x=10\) can you do the second one?

OpenStudy (anonymous):

i.e. can you solve \[x+4=0\] takes on step

OpenStudy (anonymous):

one step

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

\(x+4=0\) subtract \(4\) from both sides

OpenStudy (anonymous):

\[x+4=0\\x+4-4=0-4\\x=-4\]

OpenStudy (anonymous):

ok so its -4

OpenStudy (anonymous):

yes, so you get two answers \(x=10\) or \(x=-4\)

OpenStudy (anonymous):

-4,10?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

solve (3x+5)(2x-7)=0 @satellite73 help please last one im almost done /"

OpenStudy (anonymous):

ok here we go

OpenStudy (anonymous):

set \(3x+5=0\) first and solve for \(x\) subtract \(5\) from both sides and get \[3x=-5\] then what?

OpenStudy (anonymous):

if it is not clear let me know

OpenStudy (anonymous):

ok i got that part

OpenStudy (anonymous):

then the other side would be 2x=7

OpenStudy (anonymous):

i mean is it clear that if \(3x=-5\) then \(x=-\frac{5}{3}\) ?

OpenStudy (anonymous):

so it would be -5/3, 7/2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you done?

OpenStudy (anonymous):

i failed it

OpenStudy (anonymous):

i got a 70 and was supposed to get a 80

OpenStudy (anonymous):

damn

OpenStudy (anonymous):

can you do it again? i am fairy sure the problems we did together were right in fact i am positive

OpenStudy (anonymous):

i know they were all right but the ones i did alone D:

OpenStudy (anonymous):

damn can you try again tomorrow?

OpenStudy (anonymous):

i cant it was a test thingy now i have to do the practices

OpenStudy (anonymous):

wow good luck, i bet you can do them now in any case

OpenStudy (anonymous):

yeah thank you for your help you are very sweet C: